# Data-Driven Greenhouse Climate Regulation in Lettuce Cultivation Using BiLSTM and GRU Predictive Control

### v1

@@ Abstract
Efficient greenhouse management is essential for sustainable food production in response to a growing global population. However, maintaining optimal indoor climates requires significant energy and resources, making advanced control systems critical for economic viability and environmental sustainability. Traditional greenhouse models are often complex and imprecise, limiting the effectiveness of conventional control strategies. To address these challenges, this study investigates data-driven predictive control methods using Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) neural networks. Our experiments showed that GRU-based predictive control reduced temperature and humidity violations by up to 5% and required 40% less computation time than the LSTM approach, all while maintaining equivalent economic performance and crop yield. These findings demonstrate that GRU-based predictive control offers a more efficient and practical solution for real-time greenhouse climate regulation in precision agriculture.
@@ 1 INTRODUCTION
Global food security faces significant challenges. The number of undernourished people has been steadily increasing since 2014, reaching an estimated 720 to 811 million in 2020 [6]. This rise occurs as the world population, currently around 7.7 billion, is projected to reach 9.7 billion by 2050, driving a rapid increase in demand for fresh, nutritious food [4]. Consequently, food production must increase by approximately 70% between 2005 and 2050 to meet future needs [5]. At the same time, critical resources such as freshwater [7], fossil fuels, and arable land are becoming increasingly scarce, while the impacts of climate change continue to intensify. As a result, producing high-quality, fresh food in a resource-efficient manner is more important than ever. Greenhouse cultivation has emerged as a powerful approach to agricultural intensification. Studies indicate that it can achieve substantial yield increases, typically ranging from 10% to 20%, while simultaneously reducing critical resource inputs by 25% to 35% [8, 10]. These considerable benefits have contributed to the accelerated global growth of protected horticulture. Nevertheless, the widespread adoption of greenhouse technologies is constrained by inherent challenges, notably their high energy demands and the significant initial capital investments required [9].
@@
Greenhouses are essential for crop cultivation, as they provide a controlled environment that shields plants from unpredictable outdoor weather. In today’s world where energy costs are rising and labor and resources are becoming increasingly scarce, efficient and automated greenhouse management is more important than ever. This approach helps maximize crop production while minimizing resource use. However, outdoor weather constitutes a significant disturbance, making it challenging to control the greenhouse climate, develop accurate models, and optimize operations for optimal performance [3]. A greenhouse is a complex system consisting of two primary components: the cultivated plants and the surrounding controlled environment. This internal environment is influenced by a combination of external weather conditions, the functioning of greenhouse equipment, and the physiological activities of the plants themselves. To understand, predict, and optimize the behavior of such systems, researchers commonly rely on mathematical modeling. These models serve a dual purpose: they offer insights into the underlying physical and biological processes and support the development of control strategies to enhance system performance [1, 2].
@@
Traditional greenhouse modeling approaches are typically deterministic and based on first-principles, drawing from established scientific understanding of plant physiology, thermodynamics, and fluid dynamics [34, 33]. Although these models can provide detailed and mechanistic insights, they are often complex, featuring nonlinear dynamics and a large number of parameters that have to be manually tuned. This complexity reduces their adaptability to different greenhouse setups or operational conditions. Furthermore, first-principles models are generally limited in their ability to handle uncertainty, as estimating the variability in model parameters requires extensive datasets and computationally demanding statistical methods. While these models use data for calibration, they are not data-driven in the machine learning sense—lacking flexibility to adapt to unseen conditions without manual tuning. Building upon these modeling approaches, automated control systems have been increasingly adopted in greenhouse operations to maintain optimal environmental conditions [13]. Among advanced techniques, Model Predictive Control (MPC) stands out for its superior performance and energy efficiency relative to traditional methods [12]. The prediction model forms the cornerstone of any MPC strategy; its accuracy directly influences the quality of future state predictions and, consequently, the effectiveness of control actions. MPC typically employs analytical models—often in state-space form—to forecast system dynamics and determine optimal control inputs over a specified prediction horizon [11]. However, in practical applications, the development of first-principles models is frequently hindered by challenges such as incomplete system knowledge, complex parameter identification, and computational burdens that are incompatible with real-time operation. This has led to the emergence of data-driven black-box models as a viable alternative within the MPC framework. These models, trained exclusively on system input and output data, offer enhanced adaptability and computational efficiency. However, existing machine learning applications in greenhouse climate control, while promising, often struggle with key limitations, including poor generalization to unseen weather, sensitivity to noisy sensor data, and computational demands that impede real-time use. This research directly addresses this gap by integrating and systematically evaluating LSTM and GRU black-box models within the MPC framework for greenhouse control. The study focuses on assessing their predictive accuracy, robustness against environmental disturbances, and practical feasibility for real-time implementation under typical operational constraints and varying prediction horizons.
@@
The rest of the paper is organized as follows: Section II presents a detailed review of related work and current trends in data-driven modeling and control for greenhouse and similar systems. Section III introduces the greenhouse system under study, detailing the dataset and key environmental variables used. Section IV explains the architecture, training, and evaluation of the LSTM and GRU models. Section V describes the integration of these models into the MPC framework, outlining the control objectives, constraints, and implementation strategy. Section VI presents the results of simulation experiments, including model performance metrics and a comparative analysis of control effectiveness. Finally, Section VII concludes the paper by summarizing key findings and outlining future research directions.
@@ 2 Related work
The growing interest in deep learning for time series forecasting has been driven by its powerful ability to model complex nonlinear relationships and automatically extract high-level features from raw data—capabilities that traditional methods often lack [18, 19]. Deep learning models are particularly well-suited for handling large-scale, high-dimensional, and multi-modal datasets, enabling them to uncover intricate patterns and significantly enhance prediction accuracy [20]. These advantages have made deep learning increasingly attractive for greenhouse modelling. For example, [21] developed a hybrid model that integrates attention mechanisms, convolutional neural networks (CNN), and LSTM networks to accurately predict temperature, humidity, and $CO_{2}$ levels in edible mushroom greenhouses. Similarly, [22] demonstrated that GRU can achieve accurate minimum temperature predictions even with a limited set of input variables, outperforming other models.
In parallel, the rapid advancement of artificial intelligence and data-driven modeling techniques has led to new and effective control strategies. For instance, [23], proposed a robust data-driven model predictive control framework that maintains high-precision temperature regulation and delivers notable energy savings under uncertain conditions. These findings underscore the potential of data-driven approaches in greenhouse climate management and highlight the need for continued research in this direction. The advent of deep learning has catalyzed extensive research focused on combining the robust representational power of deep neural networks with the fields of system modeling and control.
The stability of LSTM networks is a critical factor when they are used as predictive models within a MPC framework. the study [14] addressed this issue by establishing sufficient conditions to ensure Input-to-State Stability (ISS) and Incremental Input-to-State Stability (ISS) for LSTM networks. Leveraging these stability properties, they designed an observer with guaranteed convergence. Their study demonstrated that the resulting closed-loop system combining the observer and the MPC achieves asymptotic stability.
Several studies have explored the integration of neural networks into control frameworks to reduce the computational burden of Nonlinear MPC. In this regard, [15] proposed a method that leverages artificial neural networks to learn an affine parametrization of output predictions from data. This approach enables the efficient formulation and solution of Linear Time-Varying MPC (LTV-MPC) problems for nonlinear systems. Numerical simulations on the two-tank benchmark demonstrated that the method effectively captured system dynamics and delivered strong closed-loop performance with low computational cost. Similarly, [16] showed that stable GRU can be trained and incorporated into a Nonlinear MPC framework to achieve offset-free tracking of constant references with guaranteed closed-loop stability. In their approach, a GRU was trained to identify an unknown plant while ensuring stability, based on the conditions defined in [17]. The model was further augmented with an integrator to ensure offset-free tracking, and a converging state observer was designed for the augmented system. The proposed strategy demonstrated promising results when applied to a pH neutralization process benchmark.
The study [24] proposed a reduced-dimension optimization technique that integrates recurrent neural networks into nonlinear MPC, significantly lowering computational requirements and demonstrating real-time capability in chemical reactor simulations using parallel computing. Similarly, [25] presented an innovative MPC method combining online advanced trajectory linearization with LSTM and GRU predictors, enabling simple quadratic optimization while achieving prediction accuracy and control performance comparable to fully nonlinear MPC performance in real time. [26] developed a data-driven MPC framework for synchronous motor current control using an LSTM model; due to real-time constraints, the MPC computations were done offline, and a separate neural network was trained to efficiently execute control in real time. [27] proposed a robust MPC strategy leveraging LSTM-based predictions of a preceding vehicle’s speed and position to optimize fuel consumption while accounting for prediction errors. [28] combined physics-based and data-driven modeling within a real-time MPC system employing LSTMs for hydroelectric unit control, achieving improved accuracy and reduced optimization time compared to traditional methods. Lastly, [29] introduced a Practical Nonlinear MPC approach using LSTM networks for online nonlinear system prediction and local linearization, validated through simulations on neutralization reactors.
To conclude, the existing literature underscores the growing potential of deep learning—particularly LSTM and GRU networks—for both greenhouse climate prediction and integration within MPC frameworks. These methods have demonstrated promising results across various domains by enhancing prediction accuracy, adaptability, and computational efficiency compared to traditional approaches. However, a significant gap remains in the systematic evaluation of deep recurrent models for real-time greenhouse climate control, especially under practical operational constraints and varying environmental disturbances. The novelty of this study lies in the design, integration, and evaluation of deep recurrent neural networks (LSTM and GRU) within a unified, real-time-capable MPC framework tailored to greenhouse systems. Unlike prior works that often rely on offline optimization, domain-specific configurations, or partial performance metrics, this research contributes:
  • 1.
    A comprehensive performance assessment of deep RNN-based models in MPC with respect to prediction accuracy, computational efficiency, robustness to external disturbances, and adaptability across different prediction horizons;
  • 2.
    An experimental workflow that emphasizes real-time implementation feasibility, using greenhouse-relevant input–output data to bridge the gap between theoretical model development and practical control deployment;
  • 3.
    A scalable modeling approach that minimizes reliance on system-specific tuning or first-principles knowledge, thereby enhancing transferability across greenhouse settings.
This study contributes significantly to combining deep learning with optimal control for protected agriculture, ultimately leading to more effective and resilient greenhouse management systems.

### v2

@@ Abstract
Efficient greenhouse management is essential for sustainable food production, but the high energy demand for climate regulation poses significant economic and environmental challenges. While traditional process-based greenhouse models exist, they are often too complex or imprecise for reliable control. To address this, our study introduces a novel data-driven predictive control framework using Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) neural networks within a Model Predictive Control (MPC) architecture. Training data were generated from a validated dynamic model simulating lettuce cultivation under various environmental conditions. The LSTM and GRU networks were trained to predict future greenhouse states—including temperature, humidity, CO2 concentration, and crop dry matter—with robustness confirmed via $10$-fold cross-validation. These networks were embedded into an online MPC controller to optimize heating, ventilation, and CO2 injection, aiming to minimize energy consumption and maximize crop yield while respecting biological constraints. Results showed that both the LSTM- and GRU-based controllers significantly outperformed a conventional MPC baseline. For example, humidity violations dropped from 54.77% (MPC) to 15.45% (GRU) and 17.71% (LSTM), while day–night temperature deviations were kept below $2^{\circ}\text{C}$. The GRU controller further achieved up to 40% lower computation time than its LSTM counterpart, confirming its real-time feasibility. Overall, the proposed GRU-driven predictive control approach offers a robust and computationally efficient solution for intelligent greenhouse climate automation under practical operational constraints.
@@ 1 INTRODUCTION
Global food security faces significant challenges. The number of undernourished people has increased steadily since 2014, reaching an estimated 720 to 811 million in 2020 [10]. This rise occurs as the world population, currently around 7.7 billion, is projected to reach 9.7 billion by 2050, driving a rapid increase in the demand for fresh and nutritious food [31]. Consequently, food production must increase by approximately 70% between 2005 and 2050 to meet future needs [12]. At the same time, critical resources such as freshwater [28], fossil fuels, and arable land are becoming increasingly scarce, while the impacts of climate change continue to intensify. As a result, producing high-quality fresh food in a resource-efficient manner is more important than ever. Greenhouse cultivation has emerged as a powerful approach to agricultural intensification. Studies indicate that it can achieve substantial yield increases, typically ranging from 10% to 20%, while simultaneously reducing critical resource inputs by 25% to 35% [35, 8]. These considerable benefits have contributed to the accelerated global growth of protected horticulture. Nevertheless, the widespread adoption of greenhouse technologies is limited by inherent challenges, notably their high energy demands and the significant initial capital investments required [13].
@@
Greenhouses are essential for crop cultivation, as they provide a controlled environment that shields plants from unpredictable outdoor weather. In today’s world where energy costs are rising and labor and resources are becoming increasingly scarce, efficient and automated greenhouse management is more important than ever. This approach helps maximize crop production while minimizing resource use. However, outdoor weather constitutes a significant disturbance, making it challenging to control greenhouse climate, develop accurate models, and optimize operations for optimal performance [43]. A greenhouse is a complex system consisting of two primary components: the cultivated plants and the surrounding controlled environment. This internal environment is influenced by a combination of external weather conditions, the functioning of greenhouse equipment, and the physiological activities of the plants themselves. To understand, predict and optimize the behavior of such systems, researchers commonly rely on mathematical modeling. These models serve a dual purpose: they offer information on the underlying physical and biological processes and support the development of control strategies to improve system performance [24, 19].
@@
Traditional greenhouse modeling approaches are typically deterministic and based on first-principles, drawing from established scientific understanding of plant physiology, thermodynamics, and fluid dynamics [32, 7]. Although these models can provide detailed and mechanistic insights, they are often complex, featuring nonlinear dynamics and a large number of parameters that have to be manually tuned. This complexity reduces their adaptability to different greenhouse setups or operational conditions. Furthermore, first-principles models are generally limited in their ability to handle uncertainty, as estimating the variability in model parameters requires extensive datasets and computationally demanding statistical methods. Although these models use data for calibration, they are not data-driven like machine learning models. As a result, they lack the flexibility to adapt to new conditions without manual adjustments. Building upon these modeling approaches, automated control systems have been increasingly adopted in greenhouse operations to maintain optimal environmental conditions [27]. Among advanced techniques, the MPC-controller stands out for its superior performance and energy efficiency compared to traditional methods [1]. The prediction model forms the cornerstone of any MPC strategy; its accuracy directly influences the quality of future state predictions and, consequently, the effectiveness of control actions. MPC typically employs analytical models—often in state-space form—to forecast system dynamics and determine the optimal control inputs over a specified prediction horizon [5]. However, in practical applications, the development of first-principles models is frequently hindered by challenges such as incomplete system knowledge, complex parameter identification, and computational burdens that are incompatible with real-time operation. This has led to the emergence of data-driven black-box models as a viable alternative within the MPC framework. These models, trained exclusively on system input and output data, offer enhanced adaptability and computational efficiency. However, existing machine learning applications in greenhouse climate control have shown considerable promise but still face several fundamental challenges. Models often struggle to generalize under unseen weather conditions, exhibit sensitivity to sensor noise, and require computational resources that hinder real-time deployment [16]. At the core of these issues lies a critical trade-off between maintaining expressive capacity for capturing long-term temporal dependencies and ensuring computational efficiency suitable for continuous control tasks. Although Transformer-based architectures are a dominant choice in many time-series applications, their substantial computational overhead and mixed performance evidence—where recurrent models such as GRUs and LSTMs have demonstrated superior accuracy [11, 22] in certain contexts make them less suitable for the present application.
Recurrent neural networks (RNNs)—particularly LSTM and GRU architectures—address several of these challenges through their ability to model nonlinear relationships and retain long-range temporal dependencies within sequential data. Unlike static regression models, LSTM and GRU networks can dynamically adapt to evolving system conditions, providing improved forecasting accuracy and robustness to environmental fluctuations [2]. Moreover, their compact architecture enables real-time deployment when properly optimized, making them attractive candidates for integration within MPC frameworks.
Despite these advantages, their systematic evaluation for real-time greenhouse climate control remains limited. Few studies have rigorously assessed the predictive accuracy, robustness, and computational feasibility of LSTM- and GRU-based models within operational control settings. This research aims to fill this gap by integrating and comparing LSTM and GRU black-box models within an MPC framework for greenhouse control. The study investigates their performance in terms of predictive capability, resilience to environmental disturbances, and suitability for real-time implementation under realistic operational constraints and varying prediction horizons.
@@
The remainder of this paper is organized as follows. Section II reviews related work on data-driven control approaches for greenhouse systems. Section III presents the experimental setup and dataset used in this study. Section IV introduces the architecture and performance evaluation of the proposed LSTM and GRU predictive models. Section V describes the integration of these models within a MPC framework. Section VI reports and discusses the simulation results, highlighting the control performance. Finally, Section VII concludes the paper and outlines directions for future research.
@@ 2 Related Work
Deep learning has become an increasingly prominent tool in time series forecasting due to its ability to model complex nonlinear dynamics and automatically extract hierarchical representations from raw data—capabilities that conventional statistical and mechanistic models often lack [21, 40]. Its proficiency in handling large-scale, high-dimensional, and multi-modal datasets has positioned deep neural networks as powerful alternatives for environmental and agricultural modeling tasks [23]. In the context of greenhouse systems, recurrent neural architectures such as LSTM and GRU networks have been widely investigated for their capacity to capture long-term temporal dependencies and nonlinear variable interactions. For example, [17] proposed a hybrid architecture combining attention mechanisms, convolutional layers, and LSTMs to enhance the prediction of greenhouse temperature, humidity, and $CO_{2}$ concentrations, demonstrating significant improvements in accuracy. Likewise, [15] showed that GRU networks can deliver high-fidelity temperature forecasts even when trained with limited input features, outperforming conventional data-driven methods.
Parallel advancements in artificial intelligence have catalyzed the integration of deep learning models within control frameworks, particularly MPC, to enhance greenhouse climate regulation. Data-driven MPC approaches have demonstrated their potential to improve energy efficiency and maintain precise environmental control under uncertainty [25]. The stability of recurrent models in such control settings has been examined in studies such as [38], which established sufficient conditions for Input-to-State Stability (ISS) in LSTM networks and leveraged these properties to design asymptotically stable observer-based MPC frameworks. Similarly, [6] trained stable GRU networks for plant identification within nonlinear MPC, achieving offset-free reference tracking and stability guarantees validated on benchmark processes.
Reducing the computational overhead associated with nonlinear MPC has also been a focus of research. Neural network-based surrogates have been developed to approximate system behaviors and streamline optimization processes. [29] introduced an approach using artificial neural networks to parameterize output predictions, facilitating efficient Linear Time-Varying MPC formulations for nonlinear systems. Building on this direction, [9] employed recurrent networks within MPC to lower computational requirements and demonstrated real-time capability through parallel computing architectures. Further innovations combine LSTM and GRU predictors with trajectory linearization techniques [47] or hybrid physics-informed strategies [46], achieving improved control performance and predictive precision across engineering applications, including motor drive regulation [14] and vehicle energy optimization [4].
Although these studies collectively underscore the efficacy of deep recurrent models in both prediction and control, several limitations persist. Most existing works either isolate forecasting performance or evaluate control strategies under idealized or restricted scenarios. Comprehensive investigations that assess the real-time applicability, computational feasibility, and robustness of LSTM- and GRU-based models for greenhouse climate control remain limited. Consequently, there is a distinct need for systematic evaluation frameworks that bridge prediction accuracy and closed-loop control performance under practical operational conditions. The present study addresses this gap by comparatively analyzing deep recurrent architectures against traditional modeling approaches within an integrated MPC framework for greenhouse environments, thereby extending current knowledge on data-driven strategies for sustainable climate management.